# Two Types of Gluons in QCD: Re-interpretation of ALEPH and CMS Gluon Jet Data

### v3

@@ V Re-analysis of ALEPH Gluon Jet Data by MIT-UoS Group
Figure 11: The MIT-UoS reanalysis of the ALEPH gluon jet data on the total number of the particle multiplicity. The result confirms the excess of the gluon jets with $n=1,2$, and 3.
The above ALEPF, DELPHI, and CMS gluon jet data stongly indicate the existence of two types of gluons. On the other hand, theese anormalies have not been confirmed by an independent group, so that the data need to be re-analyzed. Recently the MIT-UoS group has reanalyzed the ALEPH data. The purpose was two-fold: Is there really an anormaly in the gluon jet data? And if so, does this imply the existance of two types of gluons? Now, we discuss the outcome of the re-analysis [28].
From the 2900317 raw ALEPH $b\bar{b}g$ three jet events coming from the $Z$ boson decay obtained from the years 1991 to 1995, the MIT-UoS group selected 40984 gluon jets using the gluon tagging method similar to the one adopted by the ALEPH. With this they obtained the particle multiplicity distribution of the gluon jets for all, charged, and neutral particles. The result of the particle multiplicity for all particles is shown in Fig. 11. It has the peak at $N=6$, but also shows that the number of the events around $N=2$ and 3 is roughly 2 and 1.5 times more than the Monte Carlo prediction. This strongly indicates the existence of a second peak around $N\simeq 2.5$. This is precisely the anormaly of the particle multiplicity distribution coming from the neuron jet that we have discussed.
They also have re-analyzed the charged and neutral particle multiplicity distribution, and confirmed the existence of the same trend. For example, for the charged particle multiplicity distribution, the peak is located at $N=4$, but it has a second peak at $N=1$ which indicates the existence of the anormaly in the charged particle multiplicity distribution. With this we could conclude that the anormaly in the particle multiplicity of the ALEPH gluon jet data is real, and that we could attribute this anormaly to the existence of the neuron jets. It is really interesting that the MIT-UoS result is virtually identical to our re-interpretation of the published result of the ALEPH jet data shown in Fig. 8.
@@
With the same ALEPH data, the MIT-UoS group has re-analized the energy fragmantation of all, charged, and neutral particles coming from the gluon jets. The result of the energy fragmantation of all particles is shown in Fig. 12, which should be compared with the ALEPH result shown in Fig. 7. Here again, the result shows that the gluon jet energy fragmantation above 0.5 is 1.5 to 2 times bigger than the Monte Carlo prediction, which confirms the existence of the anormaly of the ALEPH gluon jets in large $X_{E}$ above 0.5. And obviously, this could be attributed to the existence of the neuron jets which have more energetic particles.
If this interpretation is correct, the two anormalies should be correlated, because they come from the same neuron jets. Remarkably, we can demonstrate this correlation from the ALEPH data. The MIT-UoS group have checked this correlation between all, charged, and neutral particle multiplicity and the corresponding energy fragmantation, and confirmed the existence of the correlation. The correlation for all particles is shown in Fig. 13.
It shows that 32 % of the events below $N=6$ fully expain the anormaly in the large energy fragmantation. Indeed, since this 32 % also contains the tails of the chromon jets which have high particle multiplicity, we can tells that the neuron jets which consist of 25 % of the gluon jets with the low particle multiplicity less than $N=6$ can fully account for the large energy fragmantation over 0.5 shown in Fig. 12. This confirms that there exists a strong correlation between the low particle multiplicity events and high energy fragmantation events. This is precisely the correlation predicted by the existence of the neuron jets. The MIT-UoS group also confirmed the existence of the similar correlation for the charged and neutral particles.
@@
Clearly this confirmation of the correlation between the low particle multiplicity events and high energy fragmantation events is the most compelling evidence of the exisence of the neuron jet. Of course, the excess of the low particle multiplicity shown in Fig. 8 and Fig. 11 or the excess of the high energy fragmantation shown in Fig. 7 and Fig. 12 independently implies the existence of the neuron jets. However, there is no reason why they should be correlated, unless they come from the same origin as we have predicted based on the Abelian decomposition of QCD. And the above correlation assures that indeed our interpretation is correct.
@@ VI Discussion
In this paper we have studied the five gluon jet anormalies and puzzles in the ALEPH, DELPHI, and CMS gluon jet experiments, and argued that these anormalies and puzzles could in fact be interpreted as the evidence of the existence of two types of the gluon jets. Based on the Abelian decomposition of QCD, we have argued that the excess of the low particle multiplicity and high energy fragmantation in the ALEPH data, the excess of the neutral p articles in the gluon jets in the DELPHI data, and the absence of the gluon quenching in the Pb-Pb collision in the CMS data could be interpreted as the circumferential evidence of two types of gluons in QCD.
@@
More importantly, we have discussed the existence of a new correlation between the low particle multiplicity events and high energy fragmantation events in the ALEPH data, which is obtained by the MIT-UoS group re-analyzing the ALEPH gluon jet data [28]. This correlation can be viewed a most important and persuasive evidence of the existence of the neuron jet. This is because this shows that both excesses come from the same origin, the neuron jet.
@@
Admittedly, the above evidences may not be viewed as a direct evidence of the existence of two types of gluons, since they do not identify the neuron and chromon jets separately. To have the direct evidence, we need new Monte Carlo simulations, new Pythia and FastJet predictions based on the Abelian decomposition of QCD, and test the existence of two types of gluon jets. This is beyond the scope of this paper. Nevertheless the above circumferential evidences strongly implies that indeed there could be two types of gluons in QCD. This provides us an irresistable motivation to re-analyse all existing gluon jet data more carefully.
@@
There are different ways to test the existence of two types of gluon jets. One is the jet shape. As we have pointed out, the neuron jet is expected to have a sharper jet shape, while the chromon jet has broader jet shape. So we could test the existence of the neuron jet studying the jet shape of the gluon jets. Another important feature of the neuron jet is that it has different color flow. Clearly the chromons and quarks carry color charge, but the neurons are color neutral. So the neuron jet must have different color flow. In fact, Fig. 5 suggests that the neuron jet generates an ideal color dipole pattern, but the other two jets have distorted dipole pattern. But as far as we understand, these studies have not been done by any experimental group.
@@
It should be emphasized that, in doing so we do not need any new experiment. LHC produces billions of hadron jets in a second, and ATLAS and CMS have already filed up huge data on gluon jets. Moreover, DESY, LEP, and Tevatron have old data on three jet events which could be used to confirm the existence of the neuron and chromon jets. Here again the simple number counting strongly suggests that one quarter of the gluon jets coming from the three jets events should actually be the neuron jets which do not fit to the conventional gluon jet category.
@@
There have been many experiments on the gluon jets, and one might wonder why they could not find two types of the gluon jets. There could be many reasons, but the most important reason might be that there was no theoretical motivation for them to do so, probably because they were unaware of the Abelian decomposition of QCD.
@@
But most importantly, it tells that there are two types of gluons, the color neutral neurons and colored chromons which behave differently. A common prejudice in the conventional QCD is that the eight gluons are supposed to be indistinguishable, and nobody has doubted this popular wisdom. But the Abelian decomposition tells that this is not true. In this paper we challenged this wisdom and showed how to back up this experimentally by discussing five evidences of two types of gluons. Our findings in this paper strongly implies that we may need a totally new interpretation of QCD.
At this point one might ask what are the gluon jets identified so far by experiments. Most probably they are the chromon jets, because the chromon jet has the characteristics of the gluon jet in the conventional QCD. This is evident from Fig. 5. A more interesting question is why they have not found the neuron jet. Obviously, they have not seriously searched for the neuron jet, because they had no motivation to do so. Moreover, the neuron jets consists of only 25 % of the gluon jets, and might have been systematically excluded because they do not fit to the typical gluon jet.
In fact, the sucess rate of the gluon jet tagging have been only around 70 %, which implies that the experiments have about 30% of uncertainty on the gluon jet. We believe that one reason for this uncertainty is because 25 % of the gluon jets are actually made of the neuron jets which behave differently. So it is really remarkable that the ALEPH, DELPHI, and CMS gluon jet data were able to record evidences for two types of gluon jets.
@@ Acknowledgements
The author thanks Inkyu Park and Youngkwon Jo of the MIT-UoS group for sharing their results and allowing to use some of them in this paper. The work is supported in part by National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University.

### v4

@@ V Re-analysis of ALEPH Gluon Jet Data by MIT-UoS Group
The above ALEPF, DELPHI, and CMS gluon jet data stongly indicate the existence of two types of gluons. On the other hand, these anormalies have not been confirmed by an independent group, so that the data need to be re-analyzed. Recently the MIT-UoS group has reanalyzed the ALEPH data. The purpose was two-fold: Is there really an anormaly in the gluon jet data? And if so, does this imply the existance of two types of gluons? Now, we discuss the outcome of the re-analysis [28].
Figure 12: The MIT-UoS reanalysis of the ALEPH gluon jet data on the total number of the particle multiplicity. The result confirms the excess of the gluon jet events for $n=1,2$, and 333.
From the 2900317 raw ALEPH $b\bar{b}g$ three jet events coming from the $Z$ boson decay obtained from the years 1991 to 1995, the MIT-UoS group selected 40984 gluon jets using the gluon tagging method similar to the one adopted by the ALEPH. With this they obtained the particle multiplicity distribution of the gluon jets for all, charged, and neutral particles. The result of the particle multiplicity for all particles is shown in Fig. 12. It has the peak at $N=6$, but also shows that the number of the events around $N=2$ and 3 is roughly 2 and 1.5 times more than the Monte Carlo prediction. This strongly indicates the existence of a second peak around $N\simeq 2.5$. This is precisely the anormaly of the particle multiplicity distribution coming from the nuron jet that we have discussed.
They also have re-analyzed the charged and neutral particle multiplicity distribution, and confirmed the existence of the same trend. For example, for the charged particle multiplicity distribution, the peak is located at $N=4$, but it has a second peak at $N=1$ which indicates the existence of the anormaly in the charged particle multiplicity distribution. With this we could conclude that the anormaly in the particle multiplicity of the ALEPH gluon jet data is real, and that we could attribute this anormaly to the existence of the nuron jets. It is really interesting that the MIT-UoS result is virtually identical to our re-interpretation of the published result of the ALEPH jet data shown in Fig. 8.
With the same ALEPH data, the MIT-UoS group has re-analized the energy fragmantation of all, charged, and neutral particles coming from the gluon jets. The result of the energy fragmantation of all particles is shown in Fig. 13, which should be compared with the ALEPH result shown in Fig. 7. Here again, the result shows that the gluon jet energy fragmantation above 0.5 is 1.5 to 2 times bigger than the Monte Carlo prediction, which confirms the existence of the anormaly of the ALEPH gluon jets in large $X_{E}$ above 0.5. And obviously, this could be attributed to the existence of the nuron jets which have more energetic particles.
@@
If this interpretation is correct, the two anormalies should be correlated, because they come from the same nuron jets. Remarkably, we can demonstrate this correlation from the ALEPH data. The MIT-UoS group have checked this correlation between all, charged, and neutral particle multiplicity and the corresponding energy fragmantation, and confirmed the existence of the correlation. The correlation for all particles is shown in Fig. 14.
It shows that 32 % of the events below $N=6$ fully expain the anormaly in the large energy fragmantation. Indeed, since this 32 % also contains the tails of the chromon jets which have high particle multiplicity, we can tells that the nuron jets which consist of 25 % of the gluon jets with the low particle multiplicity less than $N=6$ can fully account for the large energy fragmantation over 0.5 shown in Fig. 13. This confirms that there exists a strong correlation between the low particle multiplicity events and high energy fragmantation events. This is precisely the correlation predicted by the existence of the nuron jets. The MIT-UoS group also confirmed the existence of the similar correlation for the charged and neutral particles.
Clearly this confirmation of the correlation between the low particle multiplicity events and high energy fragmantation events is the most compelling evidence of the exisence of the nuron jet. Of course, the excess of the low particle multiplicity shown in Fig. 8 and Fig. 12 or the excess of the high energy fragmantation shown in Fig. 7 and Fig. 13 independently implies the existence of the nuron jets. However, there is no reason why they should be correlated, unless they come from the same origin as we have predicted based on the Abelian decomposition of QCD. And the above correlation assures that indeed our interpretation is correct.
@@ VI Discussion
The Abelian decomposition of QCD tells that there are two types of gluons, the color neutral nurons and colored chromons, in QCD. In this paper we have studied the four gluon jet anormalies and puzzles in the ALEPH, DELPHI, and CMS gluon jet experiments, and argued that these anormalies and puzzles could in fact be interpreted as the evidence of the existence of two types of the gluon jets. Based on the Abelian decomposition of QCD, we have argued that the excess of the low particle multiplicity and high energy fragmantation in the ALEPH data, the excess of the neutral particles in the gluon jets in the DELPHI data, and the absence of the gluon quenching in the Pb-Pb collision in the CMS data could be interpreted as the circumferential evidence of two types of gluons in QCD.
@@
More importantly, we have discussed the existence of a new correlation between the low particle multiplicity events and high energy fragmantation events in the ALEPH data, which is obtained by the MIT-UoS group re-analyzing the ALEPH gluon jet data [28]. This correlation can be viewed a most important and persuasive evidence of the existence of the nuron jet. This is because this shows that both excesses come from the same origin, the nuron jet.
@@
Admittedly, the above evidences may not be viewed as a direct evidence of the existence of two types of gluons, since they do not identify the nuron and chromon jets separately. To have the direct evidence, we need new Monte Carlo simulations, new Pythia and FastJet predictions based on the Abelian decomposition of QCD, and confirm the existence of two types of gluon jets. This is beyond the scope of this paper. Nevertheless the above circumferential evidences strongly implies that indeed there could be two types of gluons in QCD. This provides us an irresistable motivation to re-analyse all existing gluon jet data more carefully.
@@
There are different ways to test the existence of two types of gluon jets. One is the jet shape. As we have pointed out, the nuron jet is expected to have a sharper jet shape, while the chromon jet has broader jet shape. So we could test the existence of the nuron jet studying the jet shape of the gluon jets. Another important feature of the nuron jet is that it has different color flow. Clearly the chromons and quarks carry color charge, but the nurons are color neutral. So the nuron jet must have different color flow. In fact, Fig. 5 suggests that the nuron jet generates an ideal color dipole pattern, but the other two jets have distorted dipole pattern. But as far as we understand, these studies have not been done by any experimental group.
@@
It should be emphasized that, in doing so we do not need any new experiment. LHC produces billions of hadron jets in a second, and ATLAS and CMS have already filed up huge data on gluon jets. Moreover, DESY, LEP, and Tevatron have old data on three jet events which could be used to confirm the existence of the nuron and chromon jets. Here again the simple number counting strongly suggests that one quarter of the gluon jets coming from the three jets events should actually be the nuron jets which do not fit to the conventional gluon jet category.
@@
There have been many experiments on the gluon jets, and one might wonder why they could not find two types of the gluon jets. There could be many reasons, but the most important reason might be that there was no theoretical motivation for them to do so, probably because they were unaware of the possibility of two types of gluons in QCD.
@@
But most importantly, it tells that there are two types of gluons, the color neutral nurons and colored chromons which behave differently. A common prejudice in the conventional QCD is that the eight gluons are supposed to be indistinguishable, and nobody has doubted this popular wisdom. But the Abelian decomposition tells that this may not be true. In this paper we challenged this wisdom and showed how to back up this experimentally by discussing five evidences of two types of gluons. Our findings in this paper strongly implies that we may need a totally new interpretation of QCD.
At this point one might ask what are the gluon jets identified so far by experiments. Most probably they are the chromon jets, because the chromon jet has the characteristics of the gluon jet in the conventional QCD. This is evident from Fig. 5. A more interesting question is why they have not found the nuron jet. Obviously, they have not seriously searched for the nuron jet, because they had no motivation to do so. Moreover, the nuron jets consists of only 25 % of the gluon jets, and might have been systematically excluded because they do not fit to the typical gluon jet.
In fact, the sucess rate of the gluon jet tagging have been only around 70 %, which implies that the experiments have about 30% of uncertainty on the gluon jet. We believe that one reason for this uncertainty is because 25 % of the gluon jets are actually made of the nuron jets which behave differently. So it is really remarkable that the ALEPH, DELPHI, and CMS gluon jet data were able to record evidences for two types of gluon jets in spite of this uncertainty.
In this paper we have discussed the experimental verification of the Abelian decomposition. But the physical applications of the Abelian decomposition is as important as the experimental verification. Obviously a natural place to apply the Abelian decomposition is the parton distribution in QCD and have a new parton distribution function for hadrons, in particular for the proton parton distribution function. In this direction there has been an important progress. Recently the theory group of Institute of Modern Physics of Chinese Academy of Science has re-analized the Dokshitzer-Gribov-Lipatov-Altarelli-Parisi (DGLAP) evolution equation of the gluon parton distribution in proton, and suceeded to decompose it and obtain two new evolution equations for the nuron and chromon saparately [29]. This could change our understanding of the parton distribution and evolution in QCD drastically. For this to happen, of course, we first need to confirm the existence of two types of gluons in QCD experimentally. And this was precisely what we tried to do in this paper.
@@ Acknowledgements
The author thanks Yen-Jie Lee for informing the CMS data on the gluon quenching anormaly in heavy ion collision, and Inkyu Park and Youngkwon Jo for sharing the result of the MIT-UoS re-analysis of the ALEPH data and allowing to use some of them in this paper. The work is supported in part by National Research Foundation of Korea funded by the Ministry of Science and Technology (Grant 2022-R1A2C1006999) and by Center for Quantum Spacetime, Sogang University.